# Generalized Moving Least-Squares for Solving Vector-valued PDEs on Unknown Manifolds

### v1

@@ 4 GMLS using extrinsic formulation
In this section, we approximate the operators on a point cloud using the extrinsic formulation. In Section 4.1, we review the representations of the differential operators on 2D surfaces as tangential derivatives [28, 17, 13, 49, 31, 22], which could be formulated as the projection of the appropriate derivatives in the ambient space. In Section 4.2, we approximate the extrinsic formulation of the gradient and the Bochner Laplacian of vector fields using the GMLS approach [33, 14, 20, 27, 26]. The vector fields here are represented using the bases from ambient space so that the resulting discrete Bochner Laplacian is of size $3N\times 3N$ where $N$ is the number of data points and $3$ corresponds to the ambient dimension. In Section 4.3, we introduce a dimension reduction technique to convert the $3N\times 3N$ matrix to a $2N\times 2N$ matrix where $2$ corresponds to the intrinsic dimension of the surface. In Section 4.4, we derive the formula for the approximation of covariant derivatives.
@@ 4.1 Review of extrinsic formulation of differential operators on manifolds
We first review some basic concepts and define the conventional notations in extrinsic differential geometry. For any point $\mathbf{x}\in M$, the local parametrization $\iota:O\subseteq\mathbb{R}^{2}\longrightarrow M\subseteq\mathbb{R}^{3}$, is defined through the following map, $\bm{\Xi|}_{\iota^{-1}(\mathbf{x})}\bm{\ \longmapsto}\mathbf{X|}_{\mathbf{x}}$. Here, $O$ denotes a open domain that contains the point $\iota^{-1}(\mathbf{x})$, which we denoted as $\bm{\Xi}_{\iota^{-1}(\mathbf{x})}$ in the canonical coordinates $\left(\frac{\partial}{\partial\xi^{1}},\frac{\partial}{\partial\xi^{2}}\right)\Big{|}_{\iota^{-1}(\mathbf{x})}$ and $\mathbf{X|}_{\mathbf{x}}$ is the embedded point represented in the ambient coordinates $\left(\frac{\partial}{\partial X^{1}},\frac{\partial}{\partial X^{2}},\frac{\partial}{\partial X^{3}}\right)\Big{|}_{\mathbf{x}}$. The local coordinate $\left(\xi^{1},\xi^{2}\right)$ is chosen arbitrarily which can also be the same as the local Monge coordinate $(\theta_{1},\theta_{2})$ in Section 3.1. The basic idea for the extrinsic formulation is to rewrite the surface derivatives in local $\left(\xi^{1},\xi^{2}\right)$ as the projection of the derivatives in the ambient space $\left(X^{1},X^{2},X^{3}\right)$.
@@
where $\overline{\mathrm{grad}}_{\mathbb{R}^{3}}=[\partial_{X^{1}},\partial_{X^{2}},\partial_{X^{3}}]^{\top}$ is the usual gradient operator defined in $\mathbb{R}^{3}$ and $\mathrm{grad}_{g}f:=g^{ij}\frac{\partial f}{\partial\xi^{i}}\frac{\partial}{\partial\xi^{j}}$ is defined w.r.t. the Riemannian metric $g$ induced by $M$ from $\mathbb{R}^{3}$. Here, $F$ is a smooth extension of $f$ to an open subset in $\mathbb{R}^{3}$ such that $F|_{M}=f$. Note that the identity (21) holds true since for any smooth function $f\in C^{\infty}(M)$, such an extension $F$ exists and the identity does not depend on the chosen extension [30]. In the following, we will abuse the notation to use $f$ as both a function defined on a surface and its extension to an open set in $\mathbb{R}^{3}$. Let $\left\{\mathbf{e}_{1},\mathbf{e}_{2},\mathbf{e}_{3}\right\}$ be the standard orthonormal basis corresponding to the Cartesian coordinate $\left\{X^{1},X^{2},X^{3}\right\}$ in $\mathbb{R}^{3}$. Then we can rewrite (21) in component form as
@@
In a nutshell, the basic idea of the extrinsic formulation for the covariant derivative follows from the definition of the tangential connection on a submanifold $M$ of $\mathbb{R}^{n}$ (see e.g. Example 4.9 of [30]). That is, for smooth vector fields $u,y\in\mathfrak{X}\left(M\right)$, the tangential connection can be defined as
$\nabla_{u}y=\mathbf{P}\left(\bar{\nabla}_{U}Y|_{M}\right),$ (23)
@@
where $\nabla$ is the Levi-Civita connection on $M$, ¯$\bar{\nabla}:\mathfrak{X}\left(\mathbb{R}^{n}\right)\times\mathfrak{X}\left(\mathbb{R}^{n}\right)\rightarrow\mathfrak{X}\left(\mathbb{R}^{n}\right)$ is the Euclidean connection (directional derivative) on $\mathbb{R}^{n}$ mapping $(U,Y)$ to ¯$\bar{\nabla}_{U}Y$, $\mathbf{P}:T\mathbb{R}^{n}\rightarrow TM$ is the orthogonal projection onto $TM$, and $U$ and $Y$ are extensions of $u$ and $y$ onto an open subset in $\mathbb{R}^{n}$ such that $U|_{M}=u$ and $Y|_{M}=y$, respectively. One can show that such extensions exist and the value is independent of the chosen extension and thus $\nabla_{u}y$ is well defined. More details about geometry can also be found in [11, 39, 8, 3] and references therein.
@@
We now review the extrinsic formulation in our previous work [22] for the gradient and the Bochner Laplacian of a 2D vector field when it has an extension in $\mathbb{R}^{3}$. For a vector $u\in\mathfrak{X}(M),$ its gradient in the local coordinate is defined to be a $(2,0)$ tensor field in $\mathfrak{X}(M)\times\mathfrak{X}(M)$, ${\mathrm{grad}}_{g}u{=\sharp\nabla u}=g^{kj}\tilde{u}_{;k}^{i}\frac{\partial}{\partial\xi^{i}}\otimes\frac{\partial}{\partial\xi^{j}}$, where $\sharp$ is the standard musical isomorphism notation to raise the index. Let $U=U^{k}\frac{\partial}{\partial X^{k}}$ be a smooth extension of $u=u^{i}\frac{\partial}{\partial\xi^{i}}$ to an open subset in $\mathbb{R}^{3}$ satisfying $U|_{M}=u$. In fact, $\{\frac{\partial}{\partial X^{1}},\frac{\partial}{\partial X^{2}},\frac{\partial}{\partial X^{3}}\}=\left\{\mathbf{e}_{1},\mathbf{e}_{2},\mathbf{e}_{3}\right\}$ is the standard orthonormal basis for the Cartesian coordinate system. Let $U=(U^{1},U^{2},U^{3})^{\top}\in\mathbb{R}^{3\times 1}$ be the coordinate representation and let $\frac{\partial}{\partial\mathbf{X}}=(\frac{\partial}{\partial X^{1}},\frac{\partial}{\partial X^{2}},\frac{\partial}{\partial X^{3}})^{\top}\in\mathbb{R}^{3\times 1}$ be the gradient operator $\overline{{\mathrm{grad}}}_{\mathbb{R}^{3}}$. Then one can obtain the extrinsic formulation for the gradient of a vector field (see e.g. [22, 21]):
@@
${\mathrm{grad}}_{g}u$ $=$ $\mathbf{P}\left(\overline{{\mathrm{grad}}}_{\mathbb{R}^{3}}U\right)\mathbf{P}=\mathbf{P}\frac{\partial U}{\partial\mathbf{X}}\mathbf{P}=\mathbf{P}\left[\begin{matrix}\mathcal{G}_{1}U^{1}&\mathcal{G}_{2}U^{1}&\mathcal{G}_{3}U^{1}\\ \mathcal{G}_{1}U^{2}&\mathcal{G}_{2}U^{2}&\mathcal{G}_{3}U^{2}\\ \mathcal{G}_{1}U^{3}&\mathcal{G}_{2}U^{3}&\mathcal{G}_{3}U^{3}\end{matrix}\right]$
::=:= $\left(\mathcal{H}_{1}U,\mathcal{H}_{2}U,\mathcal{H}_{3}U\right),$
@@
where for any $\mathbf{x}\in M$, the projection matrix $\mathbf{P}=[P_{ij}]_{i,j=1}^{3}$ is symmetric, $\overline{{\mathrm{grad}}}_{\mathbb{R}^{n}}U=\frac{\partial U}{\partial\mathbf{X}}=[\frac{\partial U^{i}}{\partial X^{j}}]_{i,j=1}^{3}$ is a $3\times 3$ matrix, and $\mathcal{H}_{s}:=\mathbf{P}\mathrm{diag}\left(\mathcal{G}_{s},\mathcal{G}_{s},\mathcal{G}_{s}\right)$ has the form,
@@
$\mathcal{H}_{s}U:=\mathbf{P}\left[\begin{matrix}\mathcal{G}_{s}&&\\ &\mathcal{G}_{s}&\\ &&\mathcal{G}_{s}\end{matrix}\right]\left[\begin{matrix}U^{1}\\ U^{2}\\ U^{3}\end{matrix}\right]=\left[\begin{matrix}P_{11}&P_{12}&P_{13}\\ P_{21}&P_{22}&P_{23}\\ P_{31}&P_{32}&P_{33}\end{matrix}\right]\left[\begin{matrix}\mathcal{G}_{s}&&\\ &\mathcal{G}_{s}&\\ &&\mathcal{G}_{s}\end{matrix}\right]\left[\begin{matrix}U^{1}\\ U^{2}\\ U^{3}\end{matrix}\right]\in\mathbb{R}^{3\times 1}.$ (25)
@@
$\Delta_{B}u=g^{ij}u_{;ij}^{k}\frac{\partial}{\partial\xi^{k}}=\mathcal{H}_{1}\mathcal{H}_{1}U+\mathcal{H}_{2}\mathcal{H}_{2}U+\mathcal{H}_{3}\mathcal{H}_{3}U:=\bar{\Delta}_{B}U.$ (26)
Detailed derivations of the above formulation can be found in [22] or Appendix B.1.
@@ 4.2 Review of local approximation of differential operators on manifolds
$\Delta_{B}u|_{S_{{\mathbf{x}_{0}}}}$ $=$ $\bar{\Delta}_{B}U|_{S_{{\mathbf{x}_{0}}}}=\sum_{\ell=1}^{3}\mathcal{H}_{\ell}\mathcal{H}_{\ell}U|_{S_{{\mathbf{x}_{0}}}}\approx\sum_{\ell=1}^{3}\mathbf{H}_{\ell}\mathbf{H}_{\ell}\mathbf{U}_{{\mathbf{x}_{0}}}$ (30)
@@
So far, we are following most procedures in our previous work [22], where the only difference is that now we use the GMLS regression approach, which yields a sparse Laplacian matrix, instead of a global RBF interpolation in [22], which yields a dense matrix. However, the discrete Bochner Laplacian matrix is still of size $3N\times 3N$ which depends on the ambient dimension $n=3$ containing totally $9NK$ nonzero entries. For other problems with small intrinsic dimension $d$ and large ambient dimension $n$, the computation would be expensive for solving PDEs involving the Bochner Laplacian. In the next section, we introduce a dimension reduction technique to reduce the size of the Bochner Laplacian from $3N\times 3N$ to $2N\times 2N$ meanwhile reduce the number of nonzero entries from $3^{2}NK$ to $2^{2}NK$, where $2$ corresponds to the intrinsic dimension.
@@ 4.3 Reduction from ambient dimension to intrinsic dimension
For the dimension reduction, the key observation here is that the projection matrix has the decomposition as $\mathbf{P(x)=T(x)T(x)}^{\top}$ (see e.g. Proposition 2.1 of [22]) for any $\mathbf{x}\in M$, where $\mathbf{T}(\mathbf{x})=[\bm{t}_{1}(\mathbf{x}),\bm{t}_{2}(\mathbf{x})]\in\mathbb{R}^{3\times 2}$ and $\bm{t}_{i}(\mathbf{x})=(t_{i,1}(\mathbf{x}),t_{i,2}(\mathbf{x}),t_{i,3}(\mathbf{x}))^{\top}\in\mathbb{R}^{3\times 1}$ ($i=1,2$) are the same global tangent vector basis used in Section 3.2. Then, the projection matrix $\mathbf{P}^{\otimes}\in\mathbb{R}^{3K\times 3K}$ for $K$ neighboring points in $S_{{\mathbf{x}_{0}}}$ has the decomposition as $\mathbf{P}^{\otimes}=\mathbf{T}^{\otimes}\mathbf{T}^{\otimes\top},$ where $\mathbf{T}^{\otimes}\in\mathbb{R}^{3K\times 2K}$ is given by
@@
Here, $\delta_{kk}$ is defined as in (28) and $\bm{t}_{i,j}=\left(t_{i,j}(\mathbf{x}_{0,1}),\ldots,t_{i,j}(\mathbf{x}_{0,K})\right)^{\top}\in\mathbb{R}^{K\times 1}$ for $i=1,2$ and $j=1,2,3$. Then the Bochner Laplacian (30) in the stencil $S_{{\mathbf{x}_{0}}}$ can be written as
$\Delta_{B}u|_{S_{{\mathbf{x}_{0}}}}$$\approx$ $\sum_{\ell=1}^{3}\mathbf{T}^{\otimes}\left(\mathbf{T}^{\otimes\top}\left[\begin{matrix}\mathbf{G}_{\ell}&&\\ &\mathbf{G}_{\ell}&\\ &&\mathbf{G}_{\ell}\end{matrix}\right]\mathbf{T}^{\otimes}\right)\left(\mathbf{T}^{\otimes\top}\left[\begin{matrix}\mathbf{G}_{\ell}&&\\ &\mathbf{G}_{\ell}&\\ &&\mathbf{G}_{\ell}\end{matrix}\right]\mathbf{T}^{\otimes}\right)\left(\mathbf{T}^{\otimes\top}\mathbf{U}_{{\mathbf{x}_{0}}}\right)$ (32)
@@
Here, $\mathbf{R}_{\ell}:=\mathbf{T}^{\otimes\top}\left(\mathbf{I}_{3}\otimes\mathbf{G}_{\ell}\right)\mathbf{T}^{\otimes}\in\mathbb{R}^{2K\times 2K}$ with $\mathbf{I}_{3}$ being a $3\times 3$ identity matrix, $\mathbf{T}^{\otimes\top}\mathbf{U}_{{\mathbf{x}_{0}}}\in\mathbb{R}^{2K\times 1}$ is the coordinate representation of the vector field $u|_{S_{{\mathbf{x}_{0}}}}$ w.r.t. the global basis $\{\bm{t}_{j}(\mathbf{x}_{0,k})\}_{j=1,2}^{k=1,\ldots,K}$ and $\sum_{\ell=1}^{3}\mathbf{R}_{\ell}\mathbf{R}_{\ell}\in\mathbb{R}^{2K\times 2K}$ is the Bochner Laplacian matrix that maps $u|_{S_{{\mathbf{x}_{0}}}}$ to $(\Delta_{B}u)|_{S_{{\mathbf{x}_{0}}}}$ w.r.t. the global basis $\{\bm{t}_{j}(\mathbf{x}_{0,k})\}$.
In order to find the weights $\{\bm{w}^{k}\}_{k=1}^{K}$ as in (12) to approximate $\Delta_{B}u\left({\mathbf{x}_{0}}\right)$, we need to further analyze the structure of the differential matrix $\mathbf{R}_{\ell}$. From a geometric view point, the left multiplication of $\mathbf{T}^{\otimes\top}$and the right multiplication of $\mathbf{T}^{\otimes}$ in the matrix $\mathbf{R}_{\ell}=\mathbf{T}^{\otimes\top}\left(\mathbf{I}_{3}\otimes\mathbf{G}_{\ell}\right)\mathbf{T}^{\otimes}$ correspond to the pointwise left projection and the pointwise right projection of the $3K\times 3K$ matrix $\mathbf{I}_{3}\otimes\mathbf{G}_{\ell}$ onto the tangent spaces of $\{{\mathbf{x}_{0,k}}\}_{k=1}^{K}$. That is, the entries in the $s^{\text{th}}$ and ($K+s$)${}^{\text{th}}$ rows of $\mathbf{R}_{\ell}$ (w.r.t. the point $\mathbf{x}_{0,s}$) and in the $r^{\text{th}}$ and ($K+r$)${}^{\text{th}}$ columns (w.r.t. the point $\mathbf{x}_{0,r}$) can be calculated by the following projection to form a $2\times 2$ block submatrix $\mathbf{\hat{R}}_{sr}^{\ell}$,

### v2

@@ 4 GMLS using extrinsic formulation
In this section, we approximate operators on a point cloud using the extrinsic formulation. Section 4.1 reviews the representations of the differential operators on 2D surfaces as tangential derivatives [30, 18, 14, 52, 34, 24], which could be formulated as the projection of the appropriate derivatives in the ambient space. In Section 4.2, we approximate the extrinsic formulation of the gradient and the Bochner Laplacian of vector fields using the GMLS approach [37, 15, 22, 29, 28]. The vector fields here are represented using the bases from ambient space so that the resulting discrete Bochner Laplacian is of size $3N\times 3N$ where $N$ is the number of data points and $3$ corresponds to the ambient dimension. In Section 4.3, we introduce a coordinate transformation technique to convert the $3N\times 3N$ matrix to a $2N\times 2N$ matrix where $2$ corresponds to the intrinsic dimension of the surface. In Section 4.4, we directly present the generalized results of the Bochner Laplacian for $d$-dimensional manifolds in $\mathbb{R}^{n}$. In Section 4.5, we derive the approximation formula for covariant derivatives, specifically for 2D manifolds in $\mathbb{R}^{3}$.
@@ 4.1 Review of extrinsic formulation of differential operators on manifolds
We first review some basic concepts and define the conventional notations in extrinsic differential geometry for 2D manifolds in $\mathbb{R}^{3}$. For any point $\mathbf{x}\in M$, the local parametrization $\iota:O\subseteq\mathbb{R}^{2}\longrightarrow M\subseteq\mathbb{R}^{3}$, is defined through the following map, $\bm{\Xi|}_{\iota^{-1}(\mathbf{x})}\bm{\ \longmapsto}\mathbf{X|}_{\mathbf{x}}$. Here, $O$ denotes a open domain that contains the point $\iota^{-1}(\mathbf{x})$, which we denoted as $\bm{\Xi}_{\iota^{-1}(\mathbf{x})}$ in the canonical coordinates $\left(\frac{\partial}{\partial\xi^{1}},\frac{\partial}{\partial\xi^{2}}\right)\Big{|}_{\iota^{-1}(\mathbf{x})}$ and $\mathbf{X|}_{\mathbf{x}}$ is the embedded point represented in the ambient coordinates $\left(\frac{\partial}{\partial X^{1}},\frac{\partial}{\partial X^{2}},\frac{\partial}{\partial X^{3}}\right)\Big{|}_{\mathbf{x}}$. The local coordinate $\left(\xi^{1},\xi^{2}\right)$ is chosen arbitrarily which can also be the same as the local Monge coordinate $(\theta_{1},\theta_{2})$ in Section 3.1. The basic idea for the extrinsic formulation is to rewrite the surface derivatives in local $\left(\xi^{1},\xi^{2}\right)$ as the projection of the derivatives in the ambient space $\left(X^{1},X^{2},X^{3}\right)$.
@@
where $\overline{\mathrm{grad}}_{\mathbb{R}^{3}}=[\partial_{X^{1}},\partial_{X^{2}},\partial_{X^{3}}]^{\top}$ is the usual gradient operator defined in $\mathbb{R}^{3}$ and $\mathrm{grad}_{g}f:=g^{ij}\frac{\partial f}{\partial\xi^{i}}\frac{\partial}{\partial\xi^{j}}$ is defined w.r.t. the Riemannian metric $\bm{g}$ induced by $M$ from $\mathbb{R}^{3}$. Here, $F$ is a smooth extension of $f$ to an open subset in $\mathbb{R}^{3}$ such that $F|_{M}=f$. Note that the identity (33) holds true since for any smooth function $f\in C^{\infty}(M)$, such an extension $F$ exists and the identity does not depend on the chosen extension [33]. In the following, we will abuse the notation to use $f$ as both a function defined on a surface and its extension to an open set in $\mathbb{R}^{3}$. Let $\left\{\mathbf{e}_{1},\mathbf{e}_{2},\mathbf{e}_{3}\right\}$ be the standard orthonormal basis corresponding to the Cartesian coordinate $\left\{X^{1},X^{2},X^{3}\right\}$ in $\mathbb{R}^{3}$. Then we can rewrite (33) in component form as
@@
In a nutshell, the basic idea of the extrinsic formulation for the covariant derivative follows from the definition of the tangential connection on a submanifold $M$ of $\mathbb{R}^{n}$ (see e.g. Example 4.9 of [33]). That is, for smooth vector fields $\bm{u},\bm{y}\in\mathfrak{X}\left(M\right)$, the tangential connection can be defined as
$\nabla_{\bm{u}}\bm{y}=\mathbf{P}\left(\bar{\nabla}_{\bm{U}}\bm{Y}|_{M}\right),$ (35)
@@
where $\nabla$ is the Levi-Civita connection on $M$, ¯$\bar{\nabla}:\mathfrak{X}\left(\mathbb{R}^{n}\right)\times\mathfrak{X}\left(\mathbb{R}^{n}\right)\rightarrow\mathfrak{X}\left(\mathbb{R}^{n}\right)$ is the Euclidean connection (directional derivative) on $\mathbb{R}^{n}$ mapping $(\bm{U},\bm{Y})$ to ¯$\bar{\nabla}_{\bm{U}}\bm{Y}$, $\mathbf{P}:T\mathbb{R}^{n}\rightarrow TM$ is the orthogonal projection onto $TM$, and $\bm{U}$ and $\bm{Y}$ are extensions of $u$ and $y$ onto an open subset in $\mathbb{R}^{n}$ such that $\bm{U}|_{M}=\bm{u}$ and $\bm{Y}|_{M}=\bm{y}$, respectively. One can show that such extensions exist and the value is independent of the chosen extension and thus $\nabla_{\bm{u}}\bm{y}$ is well defined. More details about geometry can also be found in [11, 41, 8, 3] and references therein.
@@
We now review the extrinsic formulation in our previous work [24] for the gradient and the Bochner Laplacian of a 2D vector field when it has an extension in $\mathbb{R}^{3}$. For a vector field $\bm{u}\in\mathfrak{X}(M),$ its gradient in the local coordinate is defined to be a $(2,0)$ tensor field in $\mathfrak{X}(M)\times\mathfrak{X}(M)$, ${\mathrm{grad}}_{g}\bm{u}{=\sharp\nabla\bm{u}}=g^{kj}\tilde{u}_{;k}^{i}\frac{\partial}{\partial\xi^{i}}\otimes\frac{\partial}{\partial\xi^{j}}$, where $\sharp$ is the standard musical isomorphism notation to raise the index. Let $\bm{U}=U^{k}\frac{\partial}{\partial X^{k}}$ be a smooth extension of $\bm{u}=u^{i}\frac{\partial}{\partial\xi^{i}}$ to an open subset in $\mathbb{R}^{3}$ satisfying $\bm{U}|_{M}=\bm{u}$. In fact, $\{\frac{\partial}{\partial X^{1}},\frac{\partial}{\partial X^{2}},\frac{\partial}{\partial X^{3}}\}=\left\{\mathbf{e}_{1},\mathbf{e}_{2},\mathbf{e}_{3}\right\}$ is the standard orthonormal basis for the Cartesian coordinate system. Let $\bm{U}=(U^{1},U^{2},U^{3})^{\top}\in\mathbb{R}^{3\times 1}$ be the coordinate representation and let $\frac{\partial}{\partial\mathbf{X}}=(\frac{\partial}{\partial X^{1}},\frac{\partial}{\partial X^{2}},\frac{\partial}{\partial X^{3}})^{\top}\in\mathbb{R}^{3\times 1}$ be the gradient operator $\overline{{\mathrm{grad}}}_{\mathbb{R}^{3}}$. Then one can obtain the extrinsic formulation for the gradient of a vector field (see e.g. [24, 23]):
@@
${\mathrm{grad}}_{g}\bm{u}$ $=$ $\mathbf{P}\left(\overline{{\mathrm{grad}}}_{\mathbb{R}^{3}}\bm{U}\right)\mathbf{P}=\mathbf{P}\frac{\partial\bm{U}}{\partial\mathbf{X}}\mathbf{P}=\mathbf{P}\left[\begin{matrix}\mathcal{G}_{1}U^{1}&\mathcal{G}_{2}U^{1}&\mathcal{G}_{3}U^{1}\\ \mathcal{G}_{1}U^{2}&\mathcal{G}_{2}U^{2}&\mathcal{G}_{3}U^{2}\\ \mathcal{G}_{1}U^{3}&\mathcal{G}_{2}U^{3}&\mathcal{G}_{3}U^{3}\end{matrix}\right]$
::=:= $\left(\mathcal{H}_{1}\bm{U},\mathcal{H}_{2}\bm{U},\mathcal{H}_{3}\bm{U}\right),$
@@
where for any $\mathbf{x}\in M$, the projection matrix $\mathbf{P}=[P_{ij}]_{i,j=1}^{3}$ is symmetric, $\overline{{\mathrm{grad}}}_{\mathbb{R}^{n}}\bm{U}=\frac{\partial\bm{U}}{\partial\mathbf{X}}=[\frac{\partial U^{i}}{\partial X^{j}}]_{i,j=1}^{3}$ is a $3\times 3$ matrix, and $\mathcal{H}_{s}:=\mathbf{P}\mathrm{diag}\left(\mathcal{G}_{s},\mathcal{G}_{s},\mathcal{G}_{s}\right)$ has the form,
@@
$\mathcal{H}_{s}\bm{U}:=\mathbf{P}\left[\begin{matrix}\mathcal{G}_{s}&&\\ &\mathcal{G}_{s}&\\ &&\mathcal{G}_{s}\end{matrix}\right]\left[\begin{matrix}U^{1}\\ U^{2}\\ U^{3}\end{matrix}\right]=\left[\begin{matrix}P_{11}&P_{12}&P_{13}\\ P_{21}&P_{22}&P_{23}\\ P_{31}&P_{32}&P_{33}\end{matrix}\right]\left[\begin{matrix}\mathcal{G}_{s}&&\\ &\mathcal{G}_{s}&\\ &&\mathcal{G}_{s}\end{matrix}\right]\left[\begin{matrix}U^{1}\\ U^{2}\\ U^{3}\end{matrix}\right]\in\mathbb{R}^{3\times 1}.$ (37)
@@
$\Delta_{B}\bm{u}:=g^{ij}u_{;ij}^{k}\frac{\partial}{\partial\xi^{k}}=\mathcal{H}_{1}\mathcal{H}_{1}\bm{U}+\mathcal{H}_{2}\mathcal{H}_{2}\bm{U}+\mathcal{H}_{3}\mathcal{H}_{3}\bm{U}:=\bar{\Delta}_{B}\bm{U}.$ (38)
Detailed derivations of the above formulation can be found in [24] or Section B.1. Here, the operator $\bar{\Delta}_{B}$ with overline defined on the right hand side of (38) indicates that it involves with the extended vector field $\bm{U}$ on an open neighborhood of $M$ in $\mathbb{R}^{n}$ while one has $\bar{\Delta}_{B}\bm{U}=\Delta_{B}\bm{u}$ restricted on $M$ as shown in [24]. One can refer to Chap. 4 of [33] for more about extensions of variables and operators.
@@ 4.2 Review of local approximation of differential operators on manifolds
$(\Delta_{B}\bm{u})|_{S_{{\mathbf{x}_{0}}}}$ $=$ $(\bar{\Delta}_{B}\bm{U})|_{S_{{\mathbf{x}_{0}}}}=\sum_{\ell=1}^{3}(\mathcal{H}_{\ell}\mathcal{H}_{\ell}\bm{U})|_{S_{{\mathbf{x}_{0}}}}\approx\sum_{\ell=1}^{3}\mathbf{H}_{\ell}\mathbf{H}_{\ell}\mathbf{U}_{{\mathbf{x}_{0}}}$ (42)
@@
So far, we are following most procedures in our previous work [24], where the only difference is that now we use the GMLS regression approach, which yields a sparse Laplacian matrix, instead of a global RBF interpolation in [24], which yields a dense matrix. However, the discrete Bochner Laplacian matrix is still of size $3N\times 3N$ which depends on the ambient dimension $n=3$ containing $9NK$ total nonzero entries. For other problems with small intrinsic dimension $d$ and large ambient dimension $n$, the computation would be expensive for solving PDEs involving the Bochner Laplacian. In the next section, we introduce a dimension reduction technique to reduce the size of the Bochner Laplacian from $3N\times 3N$ to $2N\times 2N$, thus reducing the number of nonzero entries from $3^{2}NK$ to $2^{2}NK$, where $2$ corresponds to the intrinsic dimension.
@@ 4.3 Reduction from ambient dimension to intrinsic dimension
For the dimension reduction, the key observation here is that the projection matrix has the decomposition as $\mathbf{P(x)=T(x)T(x)}^{\top}$ (see e.g. Proposition 2.1 of [24]) for any $\mathbf{x}\in M$, where $\mathbf{T}(\mathbf{x})=[\bm{t}_{1}(\mathbf{x}),\bm{t}_{2}(\mathbf{x})]\in\mathbb{R}^{3\times 2}$ and $\bm{t}_{i}(\mathbf{x})=(t^{1}_{i}(\mathbf{x}),t^{2}_{i}(\mathbf{x}),t^{3}_{i}(\mathbf{x}))^{\top}\in\mathbb{R}^{3\times 1}$ ($i=1,2$) are the same global tangent vector basis used in Section 3.2. Then, the projection matrix $\mathbf{P}^{\otimes}\in\mathbb{R}^{3K\times 3K}$ for $K$ neighboring points in $S_{{\mathbf{x}_{0}}}$ has the decomposition as $\mathbf{P}^{\otimes}=\mathbf{T}^{\otimes}\mathbf{T}^{\otimes\top},$ where $\mathbf{T}^{\otimes}\in\mathbb{R}^{3K\times 2K}$ is given by
@@
Here, $\delta_{kk}$ is defined as in (40) and $\bm{t}_{i,j}=\left(t_{i}^{j}(\mathbf{x}_{0,1}),\ldots,t_{i}^{j}(\mathbf{x}_{0,K})\right)^{\top}\in\mathbb{R}^{K\times 1}$ is the $j$th entry of the $i$th tangent vector $\bm{t}_{i}$ for $i=1,2$ and $j=1,2,3$. Then the Bochner Laplacian (42) in the stencil $S_{{\mathbf{x}_{0}}}$ can be written as
$\Delta_{B}\bm{u}|_{S_{{\mathbf{x}_{0}}}}$$\approx$ $\sum_{\ell=1}^{3}\mathbf{T}^{\otimes}\left(\mathbf{T}^{\otimes\top}\left[\begin{matrix}\mathbf{G}_{\ell}&&\\ &\mathbf{G}_{\ell}&\\ &&\mathbf{G}_{\ell}\end{matrix}\right]\mathbf{T}^{\otimes}\right)\left(\mathbf{T}^{\otimes\top}\left[\begin{matrix}\mathbf{G}_{\ell}&&\\ &\mathbf{G}_{\ell}&\\ &&\mathbf{G}_{\ell}\end{matrix}\right]\mathbf{T}^{\otimes}\right)\left(\mathbf{T}^{\otimes\top}\mathbf{U}_{{\mathbf{x}_{0}}}\right)$ (44)
@@
Here, $\mathbf{R}_{\ell}:=\mathbf{T}^{\otimes\top}\left(\mathbf{I}_{3}\otimes\mathbf{G}_{\ell}\right)\mathbf{T}^{\otimes}\in\mathbb{R}^{2K\times 2K}$ with $\mathbf{I}_{3}$ being a $3\times 3$ identity matrix, $\mathbf{T}^{\otimes\top}\mathbf{U}_{{\mathbf{x}_{0}}}\in\mathbb{R}^{2K\times 1}$ is the coordinate representation of the vector field $\bm{u}|_{S_{{\mathbf{x}_{0}}}}$ w.r.t. the global basis $\{\bm{t}_{j}(\mathbf{x}_{0,k})\}_{j=1,2}^{k=1,\ldots,K}$ and $\sum_{\ell=1}^{3}\mathbf{R}_{\ell}\mathbf{R}_{\ell}\in\mathbb{R}^{2K\times 2K}$ is the Bochner Laplacian matrix that maps $\bm{u}|_{S_{{\mathbf{x}_{0}}}}$ to $(\Delta_{B}\bm{u})|_{S_{{\mathbf{x}_{0}}}}$ w.r.t. the global basis $\{\bm{t}_{j}(\mathbf{x}_{0,k})\}$.
In order to find the weights $\{\bm{w}^{k}\}_{k=1}^{K}$ as in (13) to approximate $\Delta_{B}\bm{u}\left({\mathbf{x}_{0}}\right)$, we need to further analyze the structure of the differential matrix $\mathbf{R}_{\ell}$. From a geometric view point, the left multiplication of $\mathbf{T}^{\otimes\top}$and the right multiplication of $\mathbf{T}^{\otimes}$ in the matrix $\mathbf{R}_{\ell}=\mathbf{T}^{\otimes\top}\left(\mathbf{I}_{3}\otimes\mathbf{G}_{\ell}\right)\mathbf{T}^{\otimes}$ correspond to the pointwise left projection and the pointwise right projection of the $3K\times 3K$ matrix $\mathbf{I}_{3}\otimes\mathbf{G}_{\ell}$ onto the tangent spaces of $\{{\mathbf{x}_{0,k}}\}_{k=1}^{K}$. That is, the entries in the $s^{\text{th}}$ and ($K+s$)${}^{\text{th}}$ rows of $\mathbf{R}_{\ell}$ (w.r.t. the point $\mathbf{x}_{0,s}$) and in the $r^{\text{th}}$ and ($K+r$)${}^{\text{th}}$ columns (w.r.t. the point $\mathbf{x}_{0,r}$) can be calculated by the following projection to form a $2\times 2$ block submatrix $\mathbf{\hat{R}}_{sr}^{\ell}$,